A man drove 220km in 3.5 hours. Part of the trip was at 60 km/h and the rest at 65 km/h. Find the time spent at a slower rate
1
Expert's answer
2013-01-16T11:19:33-0500
If part of the trip at 60 km/h was X hours
then the rest time is 3.5 - X
Total distance:
220 = 60*X + (3.5 - X)*65
220 -& 3.5*65 = 60*X - 65*X
5*X = 7.5
X = 1.5 h - part of the trip was at 60 km/h
3.5 - 1.5 = 2 h -part of the trip was at 65 km/h